# Contextualizing Recommendation Explanations with LLMs: A User Study

### v1

@@ 6.2 Limitations
First, we primarily recruited English-speaking users. Future research should include more diverse populations to strengthen the generalizability of our findings.
Second, the effectiveness of the explanations may have been influenced by the performance of the underlying LLMs. Although we used state-of-the-art models (GPT-4 Turbo and Claude 3.5 Sonnet), we should notice that future advancements in LLMs could further shape or change user perceptions.
Third, our distinction between contextualized and generic explanations may oversimplify the explanation types. Future studies should explore a broader variety of explanation strategies to better understand their impact in different contexts and on diverse user groups.
Lastly, our experimental design did not fully capture the dynamic real-world user interactions with recommender systems. User perceptions might change after consuming the recommended movies, which is not covered by this study. This highlights the necessity of conducting the user study in real scenarios.
@@ 7 Conclusion
Our findings show how and why recommendation explanations meet cognitive, affective, and utilitarian motivations under the U&G perspective. Our study confirms that LLM-generated contextualized explanations effectively address users’ cognitive needs and increase their consumption intentions. Notably, recommender systems should also avoid overwhelming users with irrelevant information that can hinder decision-making or erode trust. Future recommender system designs that integrate contextualized LLM-generated explanations can more effectively serve a wide range of user motivations, ultimately improving user satisfaction and leading to actionable engagement with the recommended content.
@@ Appendix A Recommendation Corpus Generation
4      Generate $r_{i1}$ from $m_{i}$
     $r_{ik}$: The $k$-th recommendation for the $i$-th movie
@@ 4      Generate $r_{i1}$ from $m_{i}$
6      Generate $r_{ik}$ from $\{m_{i},r_{i1},\dots,r_{i(k-1)}\}$
     Previous recommendations are incorporated to generate $r_{ik}$
@@ 6      Generate $r_{ik}$ from $\{m_{i},r_{i1},\dots,r_{i(k-1)}\}$
8     Generate $e_{ikc}$ and $e_{ikg}$ for $r_{ik}$
    $e_{ikc}$: Contextualized Explanation for $r_{ik}$
    $e_{ikg}$: Generic Explanation for $r_{ik}$
9     Save $r_{ik}$, $e_{ikc}$, and $e_{ikg}$ in $\mathcal{C}$
@@
Table 5: Recommendation corpus preparation steps. The corpus $\mathcal{C}$ contains recommendations $r_{ik}$ for each movie $m_{i}\in\mathcal{M}$, with contextualized explanations ($e_{ikc}$) and generic explanations ($e_{ikg}$) for each recommendation.
@@ Appendix B Interview Questions
Appendix C Supplementary Experiment (with Claude 3.5 Sonnet)
We replicated the experiment on a smaller U.S. sample ($N=140$) using Claude 3.5 Sonnet. The results align with our main study, showing that contextualization improves overall explanation quality, particularly in terms of persuasiveness, transparency, and trust. Moreover, contextualized explanations increase users’ consumption intentions. Additionally, the positive effect of RecSysUtility is maintained, even with another underlying LLM instead of GPT-4 Turbo. While some minor discrepancies emerged between this supplementary experiment and the main experiment, the core findings remain consistent. This supports the generalizability of our study’s conclusions. Figure 6, Figure 7, Figure 8, and Figure 9 show the user ratings of motivational needs and consumption intentions in this experiment.
Figure 6: Cognitive ratings by explanation type (shown variable: Cognitive). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
Figure 7: Affective ratings by explanation type (shown variable: Affective).
Figure 8: Utilitarian ratings by explanation type (shown variable: Utilitarian).
Figure 9: Consumption intentions by explanation type (shown variable: UserIntent). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
@@
--0.126
--0.228
@@
--0.499
Age
0.269*
0.234*
0.274*
0.145
Education
--0.006
--0.040
--0.052
0.053
RecSysUtility
0.217*
0.155
0.268
0.418*
RecSysFrequency
0.156
0.146
0.076
0.030
MovieConsumption
0.044
--0.018
--0.045
--0.148
Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$
Table 6: Mixed-effects linear regression results for the four dependent variables: Cognitive, Affective, Utilitarian, and UserIntent. Coefficients are shown with standard errors in parentheses.

### v2

@@ 6.2 Limitations
First, our findings may be limited in generalizability due to both the participant sample and the underlying models. We primarily recruited English-speaking users, which may not capture the full range of cultural and linguistic differences in user perceptions. Additionally, the effectiveness of the explanations likely depends on the capabilities of the underlying LLMs. Although we used state-of-the-art models (GPT-4 Turbo and Claude 3.5 Sonnet), future advancements may further influence how users interpret and respond to recommendations and explanations.
Second, our study may oversimplify the explanation generation process. Future research should explore various explanation strategies to better understand their effectiveness across different contexts. While we employed a typical cold-start recommendation scenario, providing additional user information, such as demographics or expressed preferences, can more efficiently reveal LLMs’ strong contextualization capabilities.
Lastly, our experimental design does not account for the longitudinal dynamics of real-world user interactions with recommender systems. While we capture users’ immediate perceptions of explanations and recommendations, these impressions may evolve as users continue to engage with the recommended content over time. Additionally, due to our control over user inputs, participants did not have the opportunity to express specific or niche movie preferences to the LLMs, which may play a significant role in shaping their perceptions. Moreover, the broader long-term effects, such as the cumulative impact on content consumption patterns and industry outcomes, remain open for future investigation.
@@ 7 Conclusion
Our findings show how and why recommendation explanations meet cognitive, affective, and utilitarian motivations under the U&G perspective. Our study confirms that LLM-generated contextualized explanations effectively address users’ cognitive and affective needs and increase their consumption intentions. Notably, recommender systems should also avoid overwhelming users with irrelevant information that can hinder decision-making or erode trust. Future recommender system designs that integrate contextualized LLM-generated explanations can more effectively serve a wide range of user motivations, ultimately improving user satisfaction and leading to actionable engagement with the recommended content.
@@ Appendix A Recommendation Corpus Generation
4      Generate $r_{i,1}$ from $m_{i}$
     $r_{i,k}$: The $k$-th recommendation for the $i$-th movie
@@ 4      Generate $r_{i,1}$ from $m_{i}$
6      Generate $r_{i,k}$ from $\{m_{i},r_{i,1},\dots,r_{i,(k-1)}\}$
     Previous recommendations are incorporated to generate $r_{i,k}$
@@ 6      Generate $r_{i,k}$ from $\{m_{i},r_{i,1},\dots,r_{i,(k-1)}\}$
8     Generate $e_{i,k,c}$ and $e_{i,k,g}$ for $r_{i,k}$
    $e_{i,k,c}$: Contextualized Explanation for $r_{i,k}$
    $e_{i,k,g}$: Generic Explanation for $r_{i,k}$
9     Save $r_{i,k}$, $e_{i,k,c}$, and $e_{i,k,g}$ in $\mathcal{C}$
@@
Table 5: Recommendation corpus preparation steps. The corpus $\mathcal{C}$ contains recommendations $r_{i,k}$ for each movie $m_{i}\in\mathcal{M}$, with contextualized explanations ($e_{i,k,c}$) and generic explanations ($e_{i,k,g}$) for each recommendation.
Table 5 details the procedure we use to construct the recommendation corpus $\mathcal{C}$ from a given movie shortlist $\mathcal{M}$. For each movie $m_{i}\in\mathcal{M}$, the system generates a sequence of four recommendations, indexed by $k=1$ to $4$.
At each step:
  • The variable $m_{i}$ denotes a single movie from the input shortlist.
  • The index $k$ indicates the recommendation order (which, equivalently, is the number of movies in the users viewing history used to produce the recommendation).
  • The system generates $r_{i,k}$, the $k$-th recommendation for movie $m_{i}$.
For the first recommendation ($k=1$), the system conditions only on the original movie $m_{i}$. For subsequent recommendations ($k>1$), it incorporates $m_{i}$ and all previously generated recommendations ($r_{i,1},\dots,r_{i,(k-1)}$) to produce $r_{i,k}$. This iterative approach models a realistic, context-aware recommendation process.
After generating each recommendation $r_{i,k}$, the system produces two types of explanations for users in different conditions:
  • $e_{i,k,c}$: a contextualized explanation that references the earlier items in the recommendation sequence;
  • $e_{i,k,g}$: a generic explanation that relates $r_{i,k}$ to the original movie $m_{i}$ without incorporating context.
Each recommendation $r_{i,k}$ and its corresponding explanations ($e_{i,k,c}$ and $e_{i,k,g}$) are added to the final corpus $\mathcal{C}$. This procedure ensures that the corpus captures both sequential recommendation dynamics and consistency across explanation styles.
@@ Appendix B Interview Questions
Appendix C Supplementary Experiments
C.1 Clade 3.5 Sonnet
We replicated the experiment on a smaller U.S. sample (NN=140) using Claude 3.5 Sonnet. The results align with our main study, showing that contextualization improves overall explanation quality, particularly in terms of persuasiveness, transparency, and trust. Besides, contextualized explanations increase users’ consumption intentions. While some minor discrepancies emerged between this supplementary experiment and the main experiment, the core findings remain consistent. Due to the limited sample size, some of the results are insignificant. Overall, the results of the supplement experiment support the generalizability of our study’s conclusions. Figure 5 shows the user ratings of motivational needs and consumption intentions. Figure 4 presents a more detailed comparison across dimensions. Table 6 shows the regression results.
Figure 4: Radar chart of average user perception ratings in different dimensions by explanation types (Model: Claude 3.5 Sonnet).
Figure 5: Ratings by explanation type (shown variable: Cognitive, Affective, Utilitarian, and UserIntent). Significance levels between conditions are indicated: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
@@ Appendix C Supplementary Experiments
-0.022
-0.190
0.105
-0.377
Age
0.227*
0.196
0.239*
0.115
Education
-0.043
-0.079
-0.090
0.001
RecSysUtility
0.118
0.051
0.168
0.250
RecSysFrequency
0.227
0.230
0.166
0.188
MovieConsumption
0.099
0.044
-0.006
-0.092
DisplayOrder
-0.077
-0.013
-0.011
-0.119
HistoryNum
-0.001
-0.099
-0.003
-0.007
HistoryNum $\times$ MovieConsumption
0.027
0.058
0.007
0.039
HistoryNum $\times$ RecSysUtility
0.009
-0.040
-0.021
-0.006
Table 6: Mixed-effects linear regression results. Coefficients are reported with standard errors in parentheses. Significance levels: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
C.2 Movie Genre
Observation of popular movie recommenders like Netflix and Amazon Video showed that most movies recommended tended to fall into widely favored genres such as Comedy and Adventure. Based on IMDb genre labels, we annotated each recommended movie in our experimental study as either Comedy or Adventure. To examine how genre influences user perceptions and consumption intentions, we performed the following regression analysis:
$Y_{ij}$ $=\beta_{0}+\beta_{1}\emph{Treatment}_{j}+\beta_{2}\emph{Age}_{j}+\beta_{3}\emph{Gender}_{j}+\beta_{4}\emph{Education}_{j}$
$\quad+\beta_{5}\emph{MovieConsumption}_{j}+\beta_{6}\emph{RecSysFrequency}_{j}$
$\quad+\beta_{7}\emph{RecSysUtility}_{j}+\beta_{8}\emph{HistoryNum}_{j}+\beta_{9}\emph{DisplayOrder}_{j}$
$\quad+\beta_{10}\emph{MvGenre}_{ij}+\beta_{11}\emph{MvGenre}_{j}\times\emph{Treatment}_{j}+u_{0j}+\epsilon_{ij}$
where ${Y}_{ij}$ denotes the dependent variables (i.e., Cognitive, Affective, Utilitarian, and UserIntent) for the $i$-th observation of the $j$-th subject, with intercepts $\beta_{0}$, subject-level random effects $u_{0j}$, coefficients $\beta_{1}$ through $\beta_{11}$, and an error term $\epsilon_{ij}$.
Results Regression results (Table 7) show that the genre of the recommended movie significantly influenced user perceptions. Specifically, movies categorized as Comedy were associated with lower ratings across affective ($\beta$=-0.203, pp<0.05), utilitarian ($\beta$=-0.149, pp<0.05), and behavioral intention outcomes ($\beta$=-0.415, pp<0.05), compared to Adventure movies. The effect of genre was not significantly moderated by explanation type (GE or CE), as the interaction terms were non-significant across all outcomes. These results suggest that genre perceptions persist regardless of contextualization strategies and that users may perceive adventure movies as more engaging or informative in recommendation explanations.
Cognitive Affective Utilitarian UserIntent
Intercept
4.946***
5.045***
5.189***
4.879***
Treatment (GE)
0.006
-0.058
0.034
-0.268
Treatment (CE)
0.438***
0.135
0.125
0.188
Gender (Female)
0.116
0.160
0.291***
-0.183
Age
-0.075
-0.089*
0.013
-0.148*
Education
0.037
0.028
0.002
0.038
RecSysUtility
0.306***
0.298***
0.273***
0.297***
RecSysFrequency
0.018
-0.004
-0.025
0.130
MovieConsumption
0.152**
0.174***
0.174***
0.126*
DisplayOrder
0.036
0.034
0.016
0.113**
HistoryNum
0.045**
0.068***
0.042**
0.098***
MvGenre (Com)
-0.096
-0.203**
-0.149*
-0.415***
MvGenre (Com) $\times$ Treatment (GE)
0.014
@@ C.2 Movie Genre
0.077
0.096
MvGenre (Com) $\times$ Treatment (CE)
-0.051
0.068
0.095
0.206
Table 7: Mixed-effects linear regression results. Coefficients are reported with standard errors in parentheses. Significance levels: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.